# An Iconic Heavy Hitter Algorithm Made Private

### v1

@@ 5.1 Overview
with probability $1-\delta/T$, where $\gamma_{1},\gamma_{2}:\mathbb{N}\to\mathbb{R}$ are non-decreasing in $t$. Here $f_{t}(x)$ is the exact frequency of $x$ in the prefix of length $t$. Second, all frequency estimates in $\mathcal{D}$ are $\varepsilon$-DP under single observation.
For target heavy hitter capacity $k$, we choose $\tilde{k}>k$ to isolate low-count, unstable labels. We maintain, throughout the pass, a tracked set $\mathcal{T}$ of at most $\tilde{k}$ items whose current estimates $\tilde{f}_{t}(\cdot)$ are top-$\tilde{k}$. At release time $T$, we apply a single threshold
$\tau(T)\ :=\ \frac{T}{\tilde{k}}\ +\ \gamma_{2}(T)\ +\ \gamma_{1}(T),$
@@
Intuitively, if $y$ appears in $\mathcal{T}$ but not $\mathcal{T}^{\prime}$, then when $y$ last failed to be in the other’s tracked set (at some time $t<T$), its maximum true frequency is bounded by $\frac{t}{\tilde{k}}+\gamma_{2}(t)+\gamma_{1}(t)\leq\tau(T)$. Therefore, unstable labels are suppressed by the threshold, while labels that are admitted by the threshold are protected by $\varepsilon$-DP through their frequency estimates.
@@ 3 $\hat{\mathcal{T}},\hat{C}\leftarrow\varnothing$
4 $\tau\leftarrow\max\{T/k,T/\tilde{k}+\gamma_{1}(T)+\gamma_{2}(T)\}$
@@ 5.2 Privacy
Let $X,X^{\prime}$ be neighboring streams of length $T$, and let $\mathcal{T},\mathcal{T}^{\prime}$ be the tracked sets produced by Algorithm 4. Fix $y\in\mathcal{T}\setminus\mathcal{T}^{\prime}$. Let $t<T$ be the last time at which $y$ appeared in either $X^{\prime}$ or $\mathcal{T}^{\prime}$ (whichever is later). Then
$f_{T}(y)\ \leq\ \frac{t}{\tilde{k}}+\gamma_{2}(t)+\gamma_{1}(t)\ \leq\ \frac{T}{\tilde{k}}+\gamma_{2}(T)+\gamma_{1}(T)\ =\ \tau(T).$
@@
Proof.
Since $y\notin\mathcal{T}^{\prime}$ at its last relevant time $t$, every label in $\mathcal{T}^{\prime}$ had estimate at least the $\tilde{k}-$th order statistic. The maximum value of the $\tilde{k}-$th order statistic is bounded above by $\frac{t}{\tilde{k}}+\gamma_{2}(t)$: at most $\tilde{k}$ labels can have true frequency exceeding $t/\tilde{k}$, and the sketch’s lower error bound inflates estimates by no more than $\gamma_{2}(t)$. Meanwhile, $y$’s estimate can be below its exact frequency by up to $\gamma_{1}(t)$. Combining, and using the monotonicity of $\gamma_{1},\gamma_{2}$, yields the stated bound on $\tilde{f}_{T}(y)$ at release time. ∎
Lemma 13 shows that items in the symmetric difference are safely suppressed by the single threshold, isolating a stable core that is consistent across neighboring streams. This allows us to establish that, conditioned on the Error Envelope, Algorithm 5 is as private as the underlying sketch.
Theorem 3.
If the sketch $\mathcal{D}$ is $\varepsilon$-DP under single observation and incurs the following error bound for each stream arrival
$\gamma_{1}(t)\ \geq\ \tilde{f}_{t}(x_{t})-f_{t}(x_{t})\ \geq\ \gamma_{2}(t)\qquad\forall x_{t}\in X,$
with probability at least $1-\delta/(2T)$, then Algorithm 5 is $(\varepsilon,\delta)$-differentially private.
The proof is relegated to Appendix C due to its similarities to Theorem 1.
The concrete envelopes and tail probabilities depend on $\mathcal{D}$, but the wrapper is agnostic to the internal design. For example, the non-private $\mathsf{CountMinSketch}$, with depth $d=\lceil\ln(1/\beta)\rceil$ and width $2k$ has the following error bounds (Lemma 4):
$t/k\geq\tilde{f}_{t}(x_{t})-f_{t}(x_{t})\geq 0,$
with probability 1-$\beta$. Moreover, a $\mathsf{CountMinSketch}$ has sensitivity $2d$ on neighboring streams zhao2022differentially. Thus, by Lemma 1, it is sufficient to add $\mathtt{Laplace}(2d/\varepsilon)$ noise to each cell for $\varepsilon$-differential privacy. By the Laplace tail bounds, with probability $1-\delta/2$, for $\psi=2d/\varepsilon\ln(4kd/\delta)$, the additive error from a noise random variable is in the interval $\psi,\psi]$. Therefore, by setting the depth to $d\lceil\ln(2T/\delta)\rceil$, we can use $\gamma_{1}(t)=t/k+\psi$ and $\gamma_{2}(t)=-\psi$ for the Error Envelope. A similar construction can be made for $\mathsf{CountSketch}$ using Lemma 5.
6 Experiments
@@
In this paper, we have proposed novel differentially private heavy hitters algorithms, designed for the single observation model. The primary contribution is the privatization of the $\mathsf{SpaceSaving}$ data structure, a state-of-the-art heavy hitters algorithm for non-private streams. We anticipate that its superior performance in non-private settings to transfer to the single observation differential privacy model.
To evaluate this claim empirically, we have implemented a suite of data structures and performed a comprehensive analysis of their privacy, utility, and performance trade-offs.
@@ 6.1 Implemented Algorithms
The following sketches are implemented and evaluated for heavy hitter detection:
  • Differentially private $\mathsf{MisraGries}$ lebeda2023better. As prior work, it serves as the key baseline.
@@
  • The Error Envelope Heavy Hitters implemented with both the $\mathsf{CountMinSketch}$ and $\mathsf{CountSketch}$. Error bounds are based on Lemmas 45. Note that $\mathsf{CountSketch}$ additionally requires estimation of $F_{2}$, which can be done using the rows of the sketch.
The instantiations of differentially private $\mathsf{SpaceSaving}$, $\mathsf{CountMinSketch}$, and $\mathsf{CountSketch}$ depend on the additional oversampling parameter $\tilde{k}$, which allows isolated elements to be suppressed without impeding recall. As the oversampling parameter relies on a trade-off between memory and error, to provide a fair comparison, we initialize all data structures with same value of $\tilde{k}=2k$. More specifically, the counter-based methods (including $\mathsf{MisraGries}$) are initialized with $\tilde{k}$ counters and the linear sketches are initialized with width $2\tilde{k}$. Subsequently, the goal of each algorithm is to admit items with frequency greater than $T/k$, for stream length $T$.
6.2 Datasets and Metrics
We ran experiments using both real network traffic and generated data. The network data set was drawn from anonymized traffic traces from the CAIDA 2019 dataset [^2] https://www.caida.org/data/passive/passive_2019_dataset.xml. The dataset consists of approximately $26.6$ million total source IP addresses and around 312,000312{,}000 distinct source IPs. We generated data from a skewed distribution (Zipf), varying the skew from 1.1 to 3.1 (in order to obtain meaningful distributions that produce at least one heavy hitter per run).
For both synthetic and real data we varied two parameters: the heavy hitter threshold $k$ to and the privacy parameter $\varepsilon$. Failure probability $\delta$ was set to $0.001$ for all experiments.
Following cormode2008finding, we evaluate the algorithms along several dimensions of efficiency and accuracy:
  • Memory footprint, measured in bytes.
All implementations are written in C++ and are publicly available [^3] https://github.com/rayneholland/DPHH Experiments were conducted on a Dell Latitude 7430 running Windows 11, equipped with an Intel Core i7 processor and 32GB of RAM.
@@ 6.2 Datasets and Metrics
6.3 Synthetic Data
Varying $k$.
Figure 10(a) shows that, at a fixed $k$, the linear sketches consume significantly more memory than the counter-based summaries, roughly a threefold increase. This is a result of the large depth (d>20d\!>\!20) required needed to enforce the error envelope. In practice, shallower tables (d=5d\!=\!5) often suffice. [^4] For $\mathsf{SpaceSaving}$$\mathsf{MisraGries}$, a heap-backed variant can reduce memory further, at the expense of higher update time. Update throughput (Figure 10(b)) favors the counter-based methods by a wide margin. The sketches show only a weak dependence on $k$ due to heap-based candidate tracking. Recall against the frequency threshold (Figure 10(c)) is perfect for $\mathsf{SpaceSaving}$, improves steadily for $\mathsf{MisraGries}$ as the register grows, and degrades for the linear sketches when the envelope is tightened. Precision is uniformly $1.0$ across all four algorithms (Figure 10(d)) under our thresholding rule. For ARE, $\mathsf{MisraGries}$ exhibits the largest error, $\mathsf{SpaceSaving}$ performs best, and $\mathsf{CountSketch}$ has a consistent edge over $\mathsf{CountMinSketch}$ (Figure 10(e)).
Varying skew $\phi$.
As skew increases, update time improves modestly for all methods (Figure 10(f)), reflecting fewer candidate replacements when the head of the distribution dominates. Precision remains perfect across all four algorithms (Figure 11(a)). ARE decreases with lower skew (Figure 11(b)), as is expected in frequency estimation. As before, $\mathsf{SpaceSaving}$ and $\mathsf{CountSketch}$ achieve the lowest error. Recall (Figure 11(c)) is perfect for $\mathsf{SpaceSaving}$, improves for $\mathsf{MisraGries}$ as skew decreases, and is lowest for $\mathsf{CountSketch}$ due to its larger envelope driven by the need to estimate and bound $F_{2}$.
Varying privacy $\varepsilon$.
Precision is unaffected by privacy and remains perfect (Figure 11(d)). Privacy impacts ARE primarily at tight budgets: for very small $\varepsilon$, error is influence by the noise term. The $\mathsf{CountMinSketch}$ is most sensitive because minimum-of-rows estimation biases toward negative fluctuations. This is also compounded by requiring a large depth to secure the error envelope. In this instance the error is reduced, however, there are settings where the error is increased significantly at lower $\varepsilon$ values due to this bias.
Median-of-rows in $\mathsf{CountSketch}$ composes with noise error better and yields lower ARE (Figure 11(e)). Recall (Figure 11(f)) is essentially unchanged for the counter-based methods (with a minor dip for $\mathsf{MisraGries}$ only at very small $\varepsilon$), while the sketches benefit as $\varepsilon$ increases. This is because the envelope shrinks and heavy hitters are admitted more reliably.
Main takeaway.
Across synthetic streams, $\mathsf{SpaceSaving}$ delivers the best overall utility (perfect recall and precision, lowest ARE) and the highest throughput, while using substantially fewer bytes per $\tilde{k}$ than the linear sketches. $\mathsf{CountSketch}$ consistently outperforms $\mathsf{CountMinSketch}$ on error, particularly under privacy noise. These trends collectively indicate that our $\mathsf{DP}$ instantiations preserve the empirical strengths and performance of the underlying non-private algorithms, with $\mathsf{SpaceSaving}$ retaining its advantage across all metrics.
6.4 Real Data
@@
(c) Precision vs $k$. Privacy $\varepsilon=0.1$.
@@
Precision is near-perfect across all methods (Figure 12(a)), with $\mathsf{SpaceSaving}$ admitting a small number of false positives as the threshold is relaxed. ARE generally declines with larger $k$ (Figure 12(b)). $\mathsf{CountMinSketch}$ exhibits a modest increase in error as more lower-frequency items are admitted, while $\mathsf{CountSketch}$ and $\mathsf{SpaceSaving}$ maintain error below $0.04$ throughout. Recall trends diverge (Figure 12(c)) between algorithms. Notably, $\mathsf{MisraGries}$ benefits from larger tables, whereas both linear sketches degrade as $k$ increases, a consequence of an expanding noise envelope that tightens selection and suppresses borderline heavy hitters.
@@
Precision is unaffected and remains perfect for all data structures (Figure 12(d)). ARE is essentially insensitive to $\varepsilon$ (Figure 12(e)), as estimation error from the underlying non-private summaries dominates the total error in the $\mathsf{DP}$ variants. Recall is impacted only at very small privacy budgets (Figure 12(f)), where the suppression thresholds include an $\varepsilon^{-1}$ term. This inflation particularly affects the linear sketches, while counter-based methods are largely stable.
@@ Appendix A Related Work
Heavy hitter detection has been studied extensively in the streaming model, with a broad taxonomy spanning counter-based (Misra Gries [misra1982finding], SpaceSaving [mitzenmacher2012hierarchical]), sketch-based (Count–Min Sketch [cormode2005improved], CountSketch [charikar2002finding]), and quantile-based approaches [greenwald2001space, shrivastava2004medians]. The most comprehensive empirical study is due to Cormode and Hadjieleftheriou, who implemented the principal families under uniform conditions and reported accuracy–space–throughput trade-offs that have guided practice for the last decade [cormode2008finding]. Their results show that, for arrivals-only streams, SpaceSaving achieves high accuracy with small memory and superior throughput, while sketch-based summaries offer mergeability and negative-update support at the cost of larger constants. We extend this experimental lens to the privacy setting, measuring performance of differentially private counter- and sketch-based algorithms, and observe that the original conclusions still hold under privatization.
@@
A number of works have explored differential privacy in the single-observation model [chan2012differentially, holland2025private, lebeda2023better, pagh2022improved, zhao2022differentially]. Early attempts to privatize Misra Gries used global sensitivity arguments, which implied noise scaling with the number of counters [chan2012differentially]. Lebeda and Tětek resolved this by exploiting the fine-grained structure of neighboring Misra Gries sketches [lebeda2023better]. Their analysis reduces the effective sensitivity to $O(1)$ (per coordinate) by leveraging shared offsets between neighbouring counters. Their approach yields $(\varepsilon,\delta)$-DP with worst-case error matching the best non-streaming private histograms up to constants. Despite strong practical performance, SpaceSaving—which typically dominates MG empirically has received little attention in the privacy literature.
@@
Linear sketches (Count-Min, Count Sketch) provide mergeable summaries with rigorous probabilistic error Guarantees, and have long been used for frequency estimation and frequent item detection. Recent work shows how to privatize these sketches in the single-observation DP model. Pagh and Thorup analyzed private CountSketch, giving improved utility bounds and median-of-rows estimators that compose well with noise [pagh2022improved]. Zhao et al. developed efficient private linear-sketch implementations and demonstrated strong accuracy–space trade-offs in practice [zhao2022differentially]. These mechanisms output DP frequency estimates for queried items. Producing heavy hitters, however, generally requires scanning or iterating over the universe to identify large counts—an operational hurdle in large domains.
@@
A common workaround uses a differentially private CountSketch as a frequency oracle inside a heavy-hitter pipeline [bassily2017practical]. While effective, and improve over naive iteration, current instantiations still rely on substantial post-processing requirements to prune candidates and locate frequent items. These post-processing steps require a central server. Our work bypasses this requirement and allows More broadly, heavy hitter detection has been given a lot of attention in the local DP model [bun2019heavy, erlingsson2014rappor, qin2016heavy, wu2022asymptotically], where users send one randomized report and the server reconstructs frequent items.
Appendix B Deferred Proofs Section 4
See 10

### v2

@@ 5.1 Overview
with probability $1-\delta/T$, where $\gamma_{1},\gamma_{2}:\mathbb{N}\to\mathbb{R}$ are non-decreasing in $t$. We refer to this error bound as the Error Envelope. Here $f_{t}(x)$ is the exact frequency of $x$ in the prefix of length $t$. Second, all frequency estimates in $\mathcal{D}$ are $\varepsilon$-DP under single observation.
Similar to the $\mathsf{SpaceSaving}$ strategy, for target heavy hitter capacity $k$, we choose $\tilde{k}>k$ to isolate low-count, unstable labels. We maintain, throughout the stream, a tracked set $\mathcal{T}$ of at most $\tilde{k}$ items whose arrival estimates are top-$\tilde{k}$. At release time $T$, we apply a single threshold
$\tau(T)\ :=\ \frac{T}{\tilde{k}}\ +\ \gamma_{2}(T)\ +\ 2\gamma_{1}(T),$
@@
Let $\mathcal{T}$ and $\mathcal{T}^{\prime}$, respectively, denote the tracked item sets on neighboring $X$ and $X^{\prime}$. Intuitively, if $y$ appears in $\mathcal{T}$ but not $\mathcal{T}^{\prime}$, then when $y$ last failed to be tracked on $X^{\prime}$ (at some time $t<T$), its maximum true frequency is bounded by $\frac{t}{\tilde{k}}+\gamma_{2}(t)+\gamma_{1}(t)\leq\tau(T)+\gamma_{1}(T)$. Therefore, unstable labels are suppressed by the threshold, while labels that are admitted by the threshold are protected by $\varepsilon$-DP through their frequency estimates.
@@ 3 $\hat{\mathcal{T}},\hat{C}\leftarrow\varnothing$
4 $\tau\leftarrow\max\{T/k,T/\tilde{k}+2\gamma_{1}(T)+\gamma_{2}(T)\}$
@@ 5.2 Privacy
Let $X,X^{\prime}$ be neighboring streams of length $T$, and let $\mathcal{T},\mathcal{T}^{\prime}$ be the tracked sets produced by Algorithm 4. Fix $y\in\mathcal{T}\setminus\mathcal{T}^{\prime}$. Let t1<Tt-1<T be the last time at which $y$ appeared in either $X^{\prime}$ or $\mathcal{T}^{\prime}$ (whichever is later). Then
$f_{t}(y)\ \leq\ \frac{t}{\tilde{k}}+\gamma_{2}(t)+\gamma_{1}(t).$
@@
Proof.
At time $t$, every label in $\mathcal{T}^{\prime}$ had estimate at least the $\tilde{k}-$th order statistic. The maximum value of the $\tilde{k}-$th order statistic is bounded above by ${t}/{\tilde{k}}+\gamma_{2}(t)$: at most $\tilde{k}$ labels can have true frequency exceeding $t/\tilde{k}$, and the sketch’s lower error bound inflates estimates by no more than $\gamma_{1}(t)$. Meanwhile, $y$’s estimate can be below its exact frequency by up to $\gamma_{2}(t)$. Combining, and using the monotonicity of $\gamma_{1},\gamma_{2}$, yields the stated bound on $\tilde{f}_{T}(y)$ at release time. ∎
As $\gamma_{1}(T)$ provides an upper bound on the error of each estimate in Algorithm 5, by Lemma 11, an item $x\in\mathcal{T}\Delta\mathcal{T}^{\prime}$ in the symmetric difference has frequency estimate at most
$\hat{f}_{T}(x)\leq f_{T}(x)+\gamma_{1}(T)\leq\frac{t}{\tilde{k}}+\gamma_{2}(T)+2\gamma_{1}(T)\leq\tau.$
Therefore, unstable items are safely suppressed by the single threshold. This allows us to establish that, conditioned on the Error Envelope, Algorithm 5 is as private as the underlying sketch.
Theorem 3.
If the sketch $\mathcal{D}$ is $\varepsilon$-DP under single observation and incurs the following error bound for any $x\in X$
$\gamma_{1}(t)\ \geq\ \tilde{f}_{t}(x)-f_{t}(x)\ \geq\ \gamma_{2}(t)$
with probability at least $)1-\delta/(2(T+\tilde{k}$)), then Algorithm 5 is $(\varepsilon,\delta)$-differentially private.
The proof is relegated to Appendix D due to its similarities to Theorem 1. The concrete envelopes and tail probabilities depend on $\mathcal{D}$, but the wrapper is agnostic to the internal design. We give examples of Error Envelopes for linear sketches in Appendix B.
6 Experiments
In this paper, we have proposed novel $\mathsf{DP}$ heavy hitter algorithms tailored to the single observation streaming model. Our central contribution is a $\mathsf{DP}$ adaptation of $\mathsf{SpaceSaving}$, a state-of-the-art heavy hitters algorithm in non-private streams. We posit that the strong empirical performance of $\mathsf{SpaceSaving}$ in the non-private setting carries over to the single-observation $\mathsf{DP}$ regime.
To substantiate this claim, we implement a suite of $\mathsf{DP}$ heavy hitter algorithms and conduct an empirical evaluation, analyzing the trade-offs between privacy, utility, and computational performance.
@@ 6.1 Implemented Algorithms
The following algorithms are implemented and evaluated for heavy hitter detection:
  • Differentially private $\mathsf{MisraGries}$ Lebeda and Tetek (2023). As prior work, it serves as the key baseline.
@@
  • The Error Envelope Heavy Hitters implemented with both the $\mathsf{CountMinSketch}$ and $\mathsf{CountSketch}$. The internal Error Envelopes are discussed in Appendix B. Note that the $\mathsf{CountSketch}$ additionally requires estimation of $F_{2}$, which can be done using the rows of the sketch.
The goal of each algorithm is to admit items with frequency greater than $T/k$, for stream length $T$. The instantiations of differentially private $\mathsf{SpaceSaving}$, $\mathsf{CountMinSketch}$, and $\mathsf{CountSketch}$ depend on the additional oversampling parameter $\tilde{k}$, which, through a trade-off between memory and error, allows isolated elements to be suppressed without impeding recall. For linear sketches, to accommodate the Error Envelope, we set $\tilde{k}=4k$. For the counter-based algorithms, we set $\tilde{k}=2k$. This results in a slight modification to $\mathsf{MisraGries}$, where we increase the number of counters to $\tilde{k}$ and (similar to $\mathsf{SpaceSaving}$) admit items that exceed the maximum of $T/k$ and the suppression threshold.
@@
6.2 Datasets and Metrics
We ran experiments using both real network traffic and generated synethetic data. The network data set was drawn from anonymized traffic traces from the CAIDA 2019 dataset [^2] https://www.caida.org/data/passive/passive_2019_dataset.xml. The dataset consists of approximately $26.6$ million total source IP addresses and around 312,000312{,}000 distinct source IPs. We generated synthetic data from a skewed distribution (Zipf), varying the skew from 1.1 to 2.7 (in order to obtain meaningful distributions that produce at least one heavy hitter per run).
For both synthetic and real data we varied two parameters: the heavy hitter threshold $k$ and the privacy parameter $\varepsilon$. The privacy failure probability $\delta$ was set to $0.001$ for all experiments.
Following Cormode and Hadjieleftheriou (2008), we evaluate the algorithms along several dimensions of efficiency and accuracy:
  • Memory footprint, measured in bytes.
All implementations are written in C++ and is publicly available [^3] https://github.com/rayneholland/DPHH. Experiments were conducted on a Dell Latitude 7430 running Windows 11, equipped with an Intel Core i7 processor and 32GB of RAM.
@@ 6.2 Datasets and Metrics
(c) Recall vs $k$. Privacy $\varepsilon=0.1$.
@@ 6.3 Synthetic Data
Figure 10(a) shows that, at a fixed $k$, the linear sketches consume significantly more memory than the counter-based summaries, roughly a factor of six. This disparity stems from the large depth (d>20d\!>\!20) required needed to enforce the error envelope. The counter-based algorithms require at most 240KB of memory, underscoring their compactness. Using a heap-based implementation the memory allocation of counter-based algorithm can be increased further, albeit at the cost of a slightly increased update time.
Update throughput (Figure 10(b)) favors the counter-based methods by a wide margin. The sketches show only a weak dependence on $k$ due to heap-based candidate tracking. Recall against the frequency threshold (Figure 10(c)) is perfect for $\mathsf{SpaceSaving}$, improves steadily for $\mathsf{MisraGries}$ as the number of registers grows, and degrades for the linear sketches as the envelope is tightens. This degradation arises because linear sketches require substantially more injected noise; as $T/k$ approaches the bound imposed by the privacy-induced noise (for larger $k$), recall correspondingly declines. Precision is uniformly $1.0$ across all four algorithms (Figure 10(d)) under our thresholding rule. For ARE, $\mathsf{MisraGries}$ exhibits the largest error, $\mathsf{SpaceSaving}$ performs best, and $\mathsf{CountSketch}$ has a consistent edge over $\mathsf{CountMinSketch}$ (Figure 10(e)).
Varying skew $\phi$.
As skew increases, throughput improves modestly for all methods (Figure 10(f)), reflecting fewer candidate replacements when the head of the distribution dominates. Precision remains perfect across all four algorithms (Figure 11(a)). ARE decreases with lower skew (Figure 11(b)), as is expected in frequency estimation. As before, $\mathsf{SpaceSaving}$ and $\mathsf{CountSketch}$ achieve the lowest error. Recall (Figure 11(c)) is perfect for $\mathsf{SpaceSaving}$, improves for $\mathsf{MisraGries}$ as skew decreases, and is lowest for $\mathsf{CountSketch}$ due to its larger envelope driven by the need to estimate and bound $F_{2}$.
Varying privacy $\varepsilon$.
Precision is unaffected by privacy and remains perfect (Figure 11(d)). Privacy impacts ARE primarily, in the form of higher estimation variance, at tight budgets: for very small $\varepsilon$, error is influenced by the noise term (Figure 11(e)). The $\mathsf{CountMinSketch}$ is most sensitive because minimum-of-rows estimation biases toward negative fluctuations. This is also compounded by requiring a large depth to secure the error envelope. In Contrast, median-of-rows in $\mathsf{CountSketch}$ composes with noise error better and yields a lower ARE. Recall (Figure 11(f)) is essentially unchanged for the counter-based methods (with a minor dip for $\mathsf{MisraGries}$ only at very small $\varepsilon$), while the sketches benefit as $\varepsilon$ increases. This is because the envelope shrinks and heavy hitters are admitted more reliably.
Main takeaway.
Across synthetic streams, $\mathsf{SpaceSaving}$ delivers the best overall utility (perfect recall and precision, lowest ARE) and the highest throughput, while using substantially less memory than the linear sketches. $\mathsf{CountSketch}$ consistently outperforms $\mathsf{CountMinSketch}$ on error, particularly under privacy noise. These trends collectively indicate that our $\mathsf{DP}$ instantiations preserve the empirical strengths and performance of the underlying non-private algorithms, with $\mathsf{SpaceSaving}$ retaining its advantage across all metrics.
6.4 Real Data
Varying $k$.
Precision is near-perfect across all methods (Figure 12(a)), with $\mathsf{SpaceSaving}$ admitting a small number of false positives as the threshold is relaxed. ARE generally declines with larger $k$ (Figure 12(b)). $\mathsf{CountMinSketch}$ exhibits a modest increase in error as more lower-frequency items are admitted as heavy, while $\mathsf{CountSketch}$ and $\mathsf{SpaceSaving}$ maintain error below $0.04$ throughout. Recall trends diverge (Figure 12(c)) between algorithms. Notably, $\mathsf{MisraGries}$ benefits from larger tables, whereas both linear sketches degrade as $k$ increases, a consequence of the privacy-induced noise dominating the frequency threshold $T/k$.
@@ 6.4 Real Data
Precision is unaffected and remains perfect for all data structures (Figure 12(d)). ARE is essentially insensitive to $\varepsilon$ (Figure 12(e)), as estimation error from the underlying non-private summaries dominates the total error in the $\mathsf{DP}$ variants. Recall is impacted only at very small privacy budgets (Figure 12(f)), where the suppression thresholds for linear sketches include an $\varepsilon^{-1}$ term.
@@ Appendix A Related Work
Heavy hitter detection has been studied extensively in the streaming model, with a broad taxonomy spanning counter-based ($\mathsf{MisraGries}$ [16], $\mathsf{SpaceSaving}$ [17]), sketch-based ($\mathsf{CountMinSketch}$ [9], $\mathsf{CountSketch}$ [7]), and quantile-based approaches [13, 20]. The most comprehensive empirical study is due to Cormode and Hadjieleftheriou, who implemented the principal families under uniform conditions and reported accuracy–space–throughput trade-offs [8]. Their results show that, for insertion-only streams, $\mathsf{SpaceSaving}$ achieves highest accuracy with small memory and high throughput, while sketch-based summaries offer deletion support at the cost of more memory and less throughout. We extend this experimental lens to the privacy setting, measuring performance of differentially private counter- and sketch-based algorithms, and observe that the original conclusions still hold under privatization.
@@
A number of works have explored differential privacy in the single-observation model [6, 14, 15, 18, 23]. Early attempts to privatize $\mathsf{MisraGries}$ used global sensitivity arguments, which implied noise scaling with the number of counters [6]. Lebeda and Tětek resolved this by exploiting the fine-grained structure of neighboring $\mathsf{MisraGries}$ sketches [15]. Their analysis reduces the effective sensitivity to $O(1)$ (per coordinate) by leveraging shared offsets between neighboring counters. Their approach yields $(\varepsilon,\delta)$-DP with worst-case error matching the best non-streaming private histograms up to constants. Despite strong practical performance, $\mathsf{SpaceSaving}$ (which typically dominates $\mathsf{MisraGries}$ empirically) has received little attention in the privacy literature.
@@
Linear sketches provide frequency vector summaries with rigorous probabilistic error guarantees, and have long been used for frequency estimation and frequent item detection. Recent work shows how to privatize these sketches in the single-observation $\mathsf{DP}$ model. Pagh and Thorup analyzed private $\mathsf{CountSketch}$, giving improved utility bounds, showing that the error of the query method composes nicely with Gaussian noise [18]. Zhao et al. developed efficient private linear-sketch implementations and demonstrated strong accuracy-space trade-offs in practice [23]. These mechanisms output $\mathsf{DP}$ frequency estimates for queried items. Producing heavy hitters, however, generally requires scanning or iterating over the universe to identify large counts. This is an operational hurdle in large domains.
@@
A common workaround uses a $\mathsf{DP}$ $\mathsf{CountSketch}$ as a frequency oracle inside a heavy-hitter pipeline [2]. While effective, and improve over brute-force iteration, current instantiations still rely on substantial post-processing requirements to prune candidates and locate frequent items. These post-processing steps require a central server. Our work bypasses this requirement and allows More broadly, heavy hitter detection has been given a lot of attention in the local DP model [5, 12, 19, 22], where users send one randomized report and the server reconstructs frequent items.
Appendix B Error Envelopes for Linear Sketches
The non-private $\mathsf{CountMinSketch}$ has the following error bound.
Lemma 12 (Approximation error for $\mathsf{CountMinSketch}$ [9]).
For $x\in\mathcal{U}$ and any $\beta,\eta\in(0,1)$, the estimation error of $\mathsf{CountMinSketch}$ with $d=\lceil\log(1/\beta)\rceil$ repetitions and table size $w=\lceil 2/\eta\rceil$ satisfies,
$\mathds{P}[\hat{f}(x)-f(x)>\eta F_{1}]\leq\beta,$
where $F_{1}=\sum_{y\in\mathcal{U}}f(y)$.
Thus, by setting the width to $2\tilde{k}$, the error in $\mathsf{CountMinSketch}$ is bounded by $t/\tilde{k}$, with probability $1-\beta$. Moreover, a $\mathsf{CountMinSketch}$ has sensitivity $2d$ on neighboring streams [23]. Thus, by Lemma 1, it is sufficient to add $\mathtt{Laplace}(2d/\varepsilon)$ noise to each cell for $\varepsilon$-differential privacy. By the Laplace tail bounds, with probability $1-\delta/2$, for $\psi=2d/\varepsilon\ln(4\tilde{k}d/\delta)$, the additive error from a noise random variable is in the interval $\psi,\psi]$. Therefore, by setting the depth to $d\lceil\ln(2(T+\tilde{k})/\delta)\rceil$, we can use $\gamma_{1}(t)=t/k+\psi$ and $\gamma_{2}(t)=-\psi$ for the Error Envelope. A similar construction can be made for $\mathsf{CountSketch}$ using the following result.
Lemma 13 (Approximation error for $\mathsf{CountSketch}$ [7]).
$\Pr\left[|\hat{f}(x)-f(x)|>\eta\sqrt{F_{2}/w}\right]\leq\beta,$
where $F_{2}=\sum_{y\in[n]}f(y)^{2}$ is the second moment of the frequency vector.
Appendix C Deferred Proofs Section 4
C.1 Lemma 4 Components
See 8